# Solving Oscillator Ordinary Differential Equations via Soft-constrained Physics-informed Neural Network with Small Data

### v2

@@ 3.1 Summary
On the other hand, compared to traditional numerical discretization methods (e.g., Runge-Kutta method, FEM), PINN has many advantages such as: 1) mesh-free, 2) seamlessly merging experimental data and tolerating noise, 3) effectively and efficiently evaluating new data samples, and 4) breaking the curse of dimensionality for high-dimensional, nonlinear, or inverse problems that are either not well solved or are expensive to solve by numerical discretization methods. Thus PINN is becoming a favorable catalyst for the emerging era of Digital Twins [ZZ23, YV23, YKH+24, QLWM24, SWvDP24].
(2) We generalized the mathematical framework and computational flow of soft-constrained PINN for solving ODEs/PDEs. It is clear from the composition of the loss function that PINN is driven by both data and physical information. Compared to conventional NN, PINN adds differential equations (including initial, boundary conditions) and a priori information or knowledge (e.g., the law of conservation of energy) as constrains or regularization. The working mechanism of the soft-constrained PINN and its accuracy and efficiency were experimentally verified by solving typical (covering first- and second-order, linear and nonlinear) oscillator ODEs such as Primer, Van der Pol, and Duffing oscillators.
@@ 3.2 Future Work
Based on this work, the next step could be to continue the exploration of PINN methods (both hard-constrained and soft-constrained) for solving stiff ODEs and PDEs, as well as the exploration of sampling methods [WZT+23, YKH+24], which are critical to the accuracy and efficiency of the solution. In addition, PINN combined with domain decomposition can be extended to large problems [KKL+21], and further extended to solve integral differential equations (IDEs), fractional order differential equations (FDEs) and stochastic differential equations (SDEs) [LMMK21].
Code Availability
The codes will be available at https://github.com/mikelu-shanghai/PINNtoODEwithSmallData. Readers in need can contact us.
@@
$\bm{x}$ spatial or spatial-temporal coordinates

### v4

@@ 3.1 Summary
On the other hand, compared to traditional numerical discretization methods (e.g., Runge-Kutta method, FEM), PINN has many advantages such as: 1) mesh-free, 2) seamlessly merging experimental data and tolerating noise, 3) effectively and efficiently evaluating new data samples, and 4) breaking the curse of dimensionality for high-dimensional, nonlinear, or inverse problems that are either not well solved or are expensive to solve by numerical discretization methods. Thus PINN is becoming a favorable catalyst for the emerging era of Digital Twins [34, 35, 28, 36, 37].
(2) We generalized the mathematical framework and computational flow of soft-constrained PINN for solving ODEs/PDEs. It is clear from the composition of the loss function that PINN is driven by both data and physical information. Compared to conventional NN, PINN adds differential equations (including initial, boundary conditions) and a priori information or knowledge (e.g., the law of conservation of energy) as constrains or regularization. Following and compared to [26], the working mechanism of the soft-constrained PINN and its accuracy and efficiency were experimentally verified by solving typical (covering first- and second-order, linear and nonlinear) oscillator ODEs such as Primer, Van der Pol, and Duffing oscillators.
@@ 3.2 Future Work
Based on this work, the next step could be to continue the exploration of PINN methods (both hard-constrained and soft-constrained) for solving stiff ODEs and PDEs, as well as the exploration of sampling methods [27, 28], which are critical to the accuracy and efficiency of the solution. In addition, PINN combined with domain decomposition can be extended to large problems [8, 38], and further extended to solve integral differential equations (IDEs), fractional order differential equations (FDEs) and stochastic differential equations (SDEs) [25].
[…]
Acknowledgements The authors would like to thank Prof. Hubert Baty from University of Strasbourg for his kind instruction and discussion. We would also like to thank the anonymous reviewers for their comments and suggestions.
[…]
Funding This work was sponsored by the National Natural Science Foundation of China (grant No. 51405289), and funded by Research Programm of Shanghai Jian Qiao University (KYJF21XX202122).
[…]
Conflict of interest The authors declare no conflict of interest.
[…]
Data/Code availability The data in the paper are computationally generated from the experimental codes. The codes will be available at https://github.com/mikelu-shanghai/PINNtoODEwithSmallData. (Readers in need can contact us.)
[…]
Author contribution Conceptualization & funding acquisition, Lu K.L. and Bi Z.; methodology & software & writing—original draft preparation, Lu K.L. and Su Y.M.; resources & investigation & validation, Lu K.L., Bi Z., Su Y.M. and Qiu C.; writing—review and editing, Lu K.L., Su Y.M. and Zhang W.J.; supervision, Zhang W.J.; All authors have read and agreed to the published version of the manuscript.
@@
$\boldsymbol{x}$ spatial or spatial-temporal coordinates